# Single Domain Generalization for Crowd Counting

### 14 Mar 2024 v1

@@ Abstract
Current image-based crowd counting widely employs density map regression due to its promising results. However, the method often suffers from severe performance degradation when tested on data from unseen scenarios. To address this so-called “domain shift” problem, we investigate single domain generalization (SDG) for crowd counting. The existing SDG approaches are mainly for classification and segmentation, and can hardly be extended to our case due to its regression nature and label ambiguity (i.e., ambiguous pixel-level ground truths). We propose MPCount, a novel SDG approach effective even for narrow source distribution. Reconstructing diverse features for density map regression with a single memory bank, MPCount retains only domain-invariant representations using a content error mask and attention consistency loss. It further introduces patch-wise classification as an auxiliary task to boost the robustness of density prediction to achieve highly accurate labels. Through extensive experiments on different datasets, MPCount is shown to significantly improve counting accuracy compared to the state of the art under diverse scenarios unobserved in the training data of narrow source distribution. Code is available at https://github.com/Shimmer93/MPCount.
@@ 1 Introduction
Image-based crowd counting is to estimate the number of people (or objects) in an image. The recent mainstream counting approaches are based on density map regression, where a density map with continuous pixel value is predicted from the input image, and the total crowd count is obtained by integrating all these pixel values. In training, the ground-truth density map is generated by smoothing the point annotations given by the locations of human heads. Existing training models predominantly assume that the testing data share the same source distribution, i.e., fully supervised learning. Despite encouraging experimental results, such assumption often breaks in practice due to changes in camera position, lighting, weather condition, etc. Such deviation from the training data (i.e. source domain) to unseen testing data (i.e. target domain) is known as domain shift, which severely degrades the performance of deep learning models in real-world deployment.
To mitigate the domain shift problem, domain adaptation (DA) approaches have been proposed to transfer knowledge from the source domain to the target one, typically by fine-tuning pre-trained models with a small amount of labeled and/or a large amount of unlabeled target domain data [44, 39, 2, 10]. Though DA has been demonstrated promising, its fine-tuning process is usually laborious. Furthermore, target domain data may not be easily available in reality.
@@
A common general SDG technique is to construct domain-invariant features from different augmented versions of an image. Such approach has been shown promising for various tasks [5, 34, 13]. Independent of data augmentation, many other SDG approaches have been proposed for classification and segmentation to further boost performance. Unfortunately, most of them cannot be extended to crowd counting, a regression task without category information. Furthermore, while labels in other tasks are often precise and domain-invariant, the labels for density-based crowd counting are usually ambiguous. Such “label ambiguity” stems from the fact that the density values are generated from point annotations, and hence human heads and backgrounds may be assigned to the same value under different scenarios.
@@
The recent work of [6] achieves DG for crowd counting based on dividing the source into multiple sub-domains. While encouraging, it requires a rather broad source domain distribution. We propose, for the first time, an SDG approach for crowd counting called MPCount which is not based on sub-domain division, thereof effective and extensible to narrow source distribution. We illustrate its performance as compared with a state-of-the-art scheme in Fig. 1. MPCount achieves its superior performance with the following two novel components:
@@
  • Attention memory bank (AMB) to tackle density regression: The proposed attention memory bank (AMB) takes a pair of features encoded from an crowd image and its data-augmented version as input. To cover continuous density values for regression with a finite memory size, AMB reconstructs each feature vector as an attention over memory vectors. These memory vectors automatically learn domain-invariant representations from the pair of features with different styles but similar content (i.e. crowd densities). The content error mask (CEM) is proposed to eliminate domain-related content from input features, which is characterized by a large discrepancy between instance-normalized feature element pairs. Furthermore, the newly employed attention consistency loss (ACL) enforces the similarity of attention scores produced from input features, ensuring the consistency of memory vectors in AMB.
@@
  • Patch-wise classification (PC) to tackle label ambiguity: To address the label ambiguity caused by pixel-level ground-truth density maps, MPCount employs a novel auxiliary task called patch-wise classification (PC). In this task, each crowd image is evenly divided into fixed-size square patches, and each patch is classified into two classes, i.e., containing human heads or not. The density values in the areas classified as without head are filtered out during density regression, so that the ambiguity in continuous pixel values is compensated by coarser but more accurate patch-wise binary labels.
@@
To validate MPCount, we conduct extensive experiments on various crowd counting datasets including ShanghaiTech A & B, UCF-QNRF and JHU-Crowd++. MPCount achieves excellent performance not only on traditional inter-dataset benchmarks, but also under our newly introduced challenging settings of narrow source distribution (where only data with the same image-level labels are available). In our study, MPCount is shown to reduce counting error significantly by 21.8% on SN $\rightarrow$ FH, 18.6% on FH $\rightarrow$ SN and 18.2% on B $\rightarrow$ A as compared to the state of the art.
@@ 3.2 Scheme Overview
The overall structure of our method is illustrated in  Fig. 2. During training, photometric transformations are applied to the original image $\mathcal{I}^{ori}$ to produce an augmented version $\mathcal{I}^{aug}$. The feature extractor encodes them into features $\mathcal{F}^{ori}$ and $\mathcal{F}^{aug}$, and a content error mask (CEM) $\mathcal{M}$ is applied to them. The masked features are reconstructed into ~$\tilde{\mathcal{F}}^{ori}$ and ~$\tilde{\mathcal{F}}^{aug}$ via an attention memory bank (AMB) and passed to the density head for crowd density prediction. Meanwhile, the highest-level features encoded by the feature extractor, $\mathcal{Z}^{ori}$ and $\mathcal{Z}^{aug}$, are fed into the patch-wise classification (PC) head, and the predicted PC maps (PCMs) $\hat{\mathcal{C}}^{ori}$ and $\hat{\mathcal{C}}^{aug}$ are binarized and resized to serve as a mask to filter out areas without human heads in the estimated density maps $\hat{\mathcal{D}}^{ori}$ and $\hat{\mathcal{D}}^{aug}$. The final predictions $\mathcal{D}^{\prime ori}$ and $\mathcal{D}^{\prime aug}$ are generated as two density maps with certain areas masked out by the PCMs. The entire training process is supervised under a combination of density losses, PC losses and the attention consistency loss (ACL). The inference process solely employs $\mathcal{I}^{ori}$ as input.
@@ 4.1 Datasets
Let S $\rightarrow$ T denote the case where S is the source domain and T is the target domain. The datasets are utilized in experiments under two distinct types of settings: 1) An entire dataset is regarded as a single domain: A $\rightarrow$ B / Q, B $\rightarrow$ A / Q and Q $\rightarrow$ A / B; 2) A dataset is partitioned into subsets according to image-level labels, and one subset constitutes a single domain: SD $\leftrightarrow$ SR and SN $\leftrightarrow$ FH.
@@ 4.4 Comparison with State of the Art
In  Tab. 2, we further conduct experiments on scene-specific domains SD $\leftrightarrow$ SR and weather-specific domains SN $\leftrightarrow$ FH. BL [20], MAN [17] and DCCUS [6] are trained using their released code, while IBN [23], SW [24] and ISW [5] are adapted by ourselves. We observe that the normalization/whitening-based methods cannot achieve satisfactory results, possibly because useful information is eliminated from feature statistics. DCCUS also obtains worse results than fully supervised methods, reflecting that the clustering procedure may fail to produce meaningful sub-domains when the source domain distribution is narrow. In contrast, our MPCount still performs the best among all tested methods, demonstrating its superiority in challenging conditions with narrow domain distributions.
@@ 4.5 Ablation Studies
Parameter A $\rightarrow$ B A $\rightarrow$ Q

### 05 Apr 2024 v2

@@ Abstract
Due to its promising results, density map regression has been widely employed for image-based crowd counting. The approach, however, often suffers from severe performance degradation when tested on data from unseen scenarios, the so-called “domain shift” problem. To address the problem, we investigate in this work single domain generalization (SDG) for crowd counting. The existing SDG approaches are mainly for image classification and segmentation, and can hardly be extended to our case due to its regression nature and label ambiguity (i.e., ambiguous pixel-level ground truths). We propose MPCount, a novel effective SDG approach even for narrow source distribution. MPCount stores diverse density values for density map regression and reconstructs domain-invariant features by means of only one memory bank, a content error mask and attention consistency loss. By partitioning the image into grids, it employs patch-wise classification as an auxiliary task to mitigate label ambiguity. Through extensive experiments on different datasets, MPCount is shown to significantly improve counting accuracy compared to the state of the art under diverse scenarios unobserved in the training data characterized by narrow source distribution. Code is available at https://github.com/Shimmer93/MPCount.
@@ 1 Introduction
Image-based crowd counting is to estimate the number of people (or objects) in an image. The recent mainstream counting approach is density map regression, where a density map with continuous pixel value is predicted from the input image, and the total crowd count is obtained by summing all these pixel values. During training, the ground-truth density map is generated by smoothing point annotations given by the locations of human heads. Density map regression has been demonstrated to achieve encouraging results in crowd counting.
Existing training models predominantly assume that the testing data share the same source distribution, i.e., fully supervised learning. Such assumption, however, often breaks in practice due to changes in camera position, weather condition, etc. Such deviation from the training data (i.e. source domain) to unseen testing data (i.e. target domain) is known as domain shift, which severely degrades the performance of deep learning models in real-world deployment.
To mitigate the domain shift problem, domain adaptation (DA) approaches have been proposed to transfer knowledge from the source domain to the target one, typically by fine-tuning pre-trained models with target domain data [44, 39, 2, 10]. Despite promising results, the fine-tuning process of DA is usually laborious. Furthermore, target domain data may not be easily available in reality.
@@
A common general SDG technique is to construct domain-invariant features from different augmented versions of an image. Such approach has been shown promising for various tasks [5, 34, 13]. Independent of data augmentation, many other SDG approaches have been proposed for classification and segmentation to further boost performance. Unfortunately, most of them cannot be extended to crowd counting, a regression task without category information. Moreover, while labels in classification and segmentation tasks are precise and domain-invariant, the labels for density-based crowd counting are usually ambiguous. Such “label ambiguity” stems from the fact that the density values are generated from point annotations, and hence human heads and backgrounds may be assigned to similar value under different scenarios.
@@
The recent work of [6] achieves DG for crowd counting based on dividing the source into multiple sub-domains. While encouraging, it requires a rather broad source domain distribution. We propose, for the first time, an SDG approach for crowd counting called MPCount without sub-domain division, thereof effective and extensible to narrow source distribution. We highlight its performance as compared with a state-of-the-art scheme in Fig. 1. MPCount achieves its superior performance with the following two novel components:
@@
  • An attention memory bank (AMB) to tackle density regression: MPCount employs only one attention memory bank (AMB) that takes a pair of features encoded from an crowd image and its data-augmented version as input. To cover continuous density values for regression with a finite memory size, AMB reconstructs each feature vector as an attention over memory vectors. These memory vectors learn domain-invariant representations from the pair of features of different styles but similar content (i.e. crowd densities). We present the content error mask (CEM) to eliminate domain-related content from input features, as characterized by a large discrepancy between instance-normalized feature element pairs. Furthermore, the newly employed attention consistency loss (ACL) enforces the similarity of attention scores produced from input features, ensuring the consistency of memory vectors.
@@
  • Patch-wise classification (PC) to tackle label ambiguity: To address the label ambiguity, MPCount employs a novel auxiliary task called patch-wise classification (PC). In this task, each crowd image is evenly divided into fixed-size square patches, say $16\times 16$, classified into two classes, i.e., containing human heads or not. The density values in the areas classified as without head are filtered out during density regression, so that the pixel-level ambiguity is overcome by coarser but more accurate patch-wise binary labels.
@@
To validate MPCount, we conduct extensive experiments on various crowd counting datasets including ShanghaiTech A (SHA) & B (SHB) and JHU-Crowd++. We introduce a new challenging setting of narrow source distribution, where only images of the same category, such as “snow” (SN) and “fog/haze” (FH) in JHU-Crowd++, are used in training. We demonstrate that MPCount achieves excellent performance not only on traditional inter-dataset benchmarks, but also under our newly introduced narrow source setting. Let $\mathcal{S}\rightarrow\mathcal{T}$ denote the case where the source domain is $\mathcal{S}$ and the target domain is $\mathcal{T}$. As compared to the state of the art, MPCount is shown to reduce significantly counting error by 21.8% on SN $\rightarrow$ FH, 18.6% on FH $\rightarrow$ SN, 18.2% on SHB $\rightarrow$ SHA and 9.5% on SHA $\rightarrow$ SHB.
@@ 3.2 Scheme Overview
The overall structure of our method is illustrated in  Fig. 2. During training, photometric transformations are applied to the original image $\mathcal{I}^{ori}$ to produce an augmented version $\mathcal{I}^{aug}$. The feature extractor encodes them into features $\mathcal{F}^{ori}$ and $\mathcal{F}^{aug}$, and a content error mask (CEM) $\mathcal{M}$ is applied to them. The masked features are reconstructed into ~$\tilde{\mathcal{F}}^{ori}$ and ~$\tilde{\mathcal{F}}^{aug}$ via an attention memory bank (AMB) and passed to the density head for crowd density prediction. Meanwhile, the highest-level features encoded by the feature extractor, $\mathcal{Z}^{ori}$ and $\mathcal{Z}^{aug}$, are fed into the patch-wise classification (PC) head, and the predicted PC maps (PCMs) $\hat{\mathcal{C}}^{ori}$ and $\hat{\mathcal{C}}^{aug}$ are binarized and resized to serve as a mask to filter out areas without human heads in the estimated density maps $\hat{\mathcal{D}}^{ori}$ and $\hat{\mathcal{D}}^{aug}$. The final predictions $\mathcal{D}^{\prime ori}$ and $\mathcal{D}^{\prime aug}$ are two density maps with certain areas masked out by the PCMs. The entire training process is supervised under a combination of density losses, PC losses and the attention consistency loss (ACL). The inference process solely employs $\mathcal{I}^{ori}$ as input, and outputs a single density map $\mathcal{D}^{\prime ori}$.
@@ 4.1 Datasets
Let $\mathcal{S}\rightarrow\mathcal{T}$ denote the case where $\mathcal{S}$ is the source domain and $\mathcal{T}$ is the target domain. The datasets are utilized in experiments under two distinct types of settings: 1) An entire dataset is regarded as a single domain: A $\rightarrow$ B / Q, B $\rightarrow$ A / Q and Q $\rightarrow$ A / B; 2) A dataset is partitioned into subsets according to image-level labels, and one subset constitutes a single domain: SD $\leftrightarrow$ SR and SN $\leftrightarrow$ FH.
@@ 4.4 Comparison with State of the Art
In  Tab. 2, we further conduct experiments on scene-specific domains SD $\leftrightarrow$ SR and weather-specific domains SN $\leftrightarrow$ FH. BL [20], MAN [17], DAOT [48] and DCCUS [6] are trained using their released code, while IBN [23], SW [24] and ISW [5] are adapted by ourselves. We observe that the normalization/whitening-based methods cannot achieve satisfactory results, possibly because useful information is eliminated from feature statistics. DCCUS also obtains worse results than fully supervised methods, reflecting that the clustering procedure may fail to produce meaningful sub-domains when the source domain distribution is narrow. In contrast, our MPCount still performs the best among all tested methods, demonstrating its superiority in challenging conditions with narrow domain distributions.
@@ 4.5 Ablation Studies
Parameter Property A $\rightarrow$ B A $\rightarrow$ Q